OpenStudy (ineedhelpplz):

PLEASE HELP!!! I WILL DO ANYTHING I WILL SELL MY SOUL IDC PLEASE HELP!!!!! Jeff is designing the seating arrangement for a concert in his high school theater. To give everyone a good view, each row must have 5 more seats than the row before it, and the 1st row can only have 12 seats. Help Jeff plan the rest of the seating by solving for how many seats are in row 25. Then explain to Jeff how to create an equation to predict the number of seats in any row. Show your work, and use complete sentences. (10 points)

OpenStudy (tkhunny):

Row 1: 12 Row 2: 12 + 5 = 17 Row 3: 12 + 2*5 = 22 Row 4: 12 + 3*5 = 27 ... Row 10: 12 + 9*5 = 57 ... Row 16: 12 + 15*5 = 87 ... Thinking, ... Thinking... Are we seeing a pattern?

OpenStudy (ineedhelpplz):

Are you multiplying 5 each time because thats the rate ?

OpenStudy (mathmale):

That's observant of you. Think: Is this in any way related to arithmetic sequences? or to the exalted brother of arithmetic sequences, geometric sequences? To answer your question, yes, but I would use the words "common ratio" instead of "rate."

OpenStudy (mathmale):

I advise you not sell your soul. Nothing's worth your soul. IDC: I do care.

OpenStudy (mathmale):

Welcome to OpenStudy!

OpenStudy (ineedhelpplz):

THANK YOU ILY

OpenStudy (tkhunny):

" ... each row must have 5 more seats than the row before it..."

OpenStudy (mathmale):

That fact eliminates one of the choices I suggested. Which one will you keep?